Six times the sum of a number and 15 is 99 more than three times the number. what is the number?
Answer by checkley77(12844)

6x+15=99+3x
6x-3x=99-15
3x=84
x=84/3
x=28 ans.
Proof:
6*28+15=99+3*28
168+15=99+84
183=183